re: 0 of 1 pt 6 of 6 (5 completel )CP1-55 Continuing Problem (book/static) HW Score: 4404% 264 anyon Canoe Company is a servIce based company that rents canoes for use on local lakes and iyers Amber and Zack Whon graduated rom coege about 10 years ag Because they both love the outdoors, they decided to begin a new business that will combine their love of outdoor acsvites wth thow busess knowledge Amber and Zack decide that they wi creaa short The business began operations on November 1, 2018 (Cick the icon to view the transactions) EOueston They bo woked tr one of the Ba Foaccounting and becane copoation Canyon Cance Company or COC Read the teaurements Requirement 1. Analyze the eflects of Canyon Canoe Company's transactions on the accounting eqalon ncludod are he folowng headngs Cash Acrounts Recovable ARI Oce Spples (O Sp1 Canoes Aerounts Payable ie Pa (Ut Pay ). Telephone Payable (Tele. Pay) Common Stock (CS), Dvidends (Dw) Canoe Rental FenetRev) Ren Expse ent Exp s Expeme tExp1 Wages Epeme (Wage Exp and Tehone Expese (ele Ee Analyze the events chronologically, one ransaction at a bime, beginning with the transaction on Novemper 1 For cach tansaction that follows the Novenber 1il aacon caa e anceechaunt a equation After calculateg the ending balance of each account on Novenber 30 calate otal ts and tot s and equty Conplee only he necessary answer boes for your ancon down all balances to the "Balne including zeo balance accounts enteing for any zoro balaces decreaseaaccount whmingn or pa ses anlyng cton the accounting Assets Lisblities Cquty Tel GS Rent age Tole Cash AR Oft Cances AP Rev Expr Sup Pay Pay 75 000 so 400) 0 00 14 300 2.20- 0) - 400 0150 lC Clear Al Enter any number in the edt elds and then clcK Check Answer 17 parts remaining
kes and rivers. Amber and Zack Wilson graduated from college about 10 years ago. They both worke e that they wi - X More Info Nov. 1 Received $16,000 cash to begin the company and issued common Office Suppli Wages Expense stock to Amber and Zack 2 Signed a lease for a building and paid $1,200 for the first month's late the balanc for your transac rent 3 Purchased canoes for $4,800 on account 4 Purchased office supplies on ac 7 Eamed $1.400 cash for rental of canoes. count, $750 13 Paid $1,500 cash for wages 15 Paid $50 dividends to stockholders. 16 Received a bill for $150 for utilities. (Use separate payable account) 20 Received a bill for $175 for cell phone expenses (Use separate payable account.) 22 Rented canoes to Early Start Daycare on account $3,000 26 Paid $1.000 on account related to the November 3, 2018, purchase 1751 28 Received $750 from Early Start Daycare for canoe rental on November 22, 2018 30 Paid $100 dividends to stockholders Print Done Clear All
Answer #1
Assets = Liabilities + Equity
Cash A/R Off. Sup. Canoes = A/P Util. Pay. Tele. Pay. + CS Div. Rent Rev. Rent Exp. Util. Exp. Wage Exp. Tele. Exp.
1 16000 = + 16000
2 -1200 = + 1200
Bal. 14800 0 0 0 = 0 0 0 + 16000 0 0 1200 0 0 0
3 4800 = 4800 +
Bal. 14800 0 0 4800 = 4800 0 0 + 16000 0 0 1200 0 0 0
4 750 = 750 +
Bal. 14800 0 750 4800 = 5550 0 0 + 16000 0 0 1200 0 0 0
7 1400 = + 1400
Bal. 16200 0 750 4800 = 5550 0 0 + 16000 0 1400 1200 0 0 0
13 -1500 = + 1500
Bal. 14700 0 750 4800 = 5550 0 0 + 16000 0 1400 1200 0 1500 0
15 -50 = + 50
Bal. 14650 0 750 4800 = 5550 0 0 + 16000 50 1400 1200 0 1500 0
16 = 150 + 150
Bal. 14650 0 750 4800 = 5550 150 0 + 16000 50 1400 1200 150 1500 0
20 = 175 + 175
Bal. 14650 0 750 4800 = 5550 150 175 + 16000 50 1400 1200 150 1500 175
22 3000 = + 3000
Bal. 14650 3000 750 4800 = 5550 150 175 + 16000 50 4400 1200 150 1500 175
26 -1000 = -1000 +
Bal. 13650 3000 750 4800 = 4550 150 175 + 16000 50 4400 1200 150 1500 175
28 750 -750 = +
Bal. 14400 2250 750 4800 = 4550 150 175 + 16000 50 4400 1200 150 1500 175
30 -100 = + 100
Bal. 14300 2250 750 4800 = 4550 150 175 + 16000 150 4400 1200 150 1500 175
TOTAL ASSETS 22100 = TOTAL LIABILITIES AND EQUITY 22100